# ClassContrast: Bridging the Spatial and Contextual Gaps for Node Representations

### v1

@@ 4.3 ClassContrast Embedding
Definition 4.4 (ClassContrast Embedding).
The final embedding $\vec{\gamma}(u)$ for node $u$ is a concatenated vector of these embeddings, resulting in a $5\cdot N$-dimensional vector where $N$ is the number of classes. For clarity, we represent $\vec{\gamma}(u)$ in a 2D format ($m\times N$) where each column corresponds to one class, and each row represents one type of "contrast" vector:
$\vec{\gamma}(u)=\left[{\begin{array}[]{lcr}\longleftarrow&\vec{\alpha}_{1i}(u)&\longrightarrow\\ \longleftarrow&\vec{\alpha}_{1o}(u)&\longrightarrow\\ \longleftarrow&\vec{\alpha}_{2}(u)&\longrightarrow\\ \longleftarrow&\vec{\beta}_{1}(u)&\longrightarrow\\ \longleftarrow&\vec{\beta}_{2}(u)&\longrightarrow\end{array}}\right]{\begin{array}[]{ll}\mbox{Spatial}&\mbox{Incoming 1-ngbd}\\ \mbox{Spatial}&\mbox{Outgoing 1-ngbd}\\ \mbox{Spatial}&\mbox{2-ngbd}\\ \mbox{Contextual}&\{d(\mathcal{X}_{u},\xi^{1}_{j})\}\\ \mbox{Contextual}&\{d(\mathcal{X}_{u},\xi^{2}_{j})\}\end{array}}$
4.4 Homophily and ClassContrast
In this section, we develop the final component of the energy landscape theory for ClassContrast that will enable us to quantify the energy landscapes between classes.
In recent years, several metrics have been introduced to study the effect of homophily on graph representation learning (Lim et al. 2021b; Jin et al. 2022; Luan et al. 2022; Platonov et al. 2023) (see overview in Section B.1). A widely used metric, the node homophily ratio, is defined as $\mathcal{H}_{\mbox{node}}(\mathcal{G})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{\eta(v)}{deg(v)}$, where $\eta(v)$ represents the number of adjacent nodes to $v$ sharing the same class. A graph $\mathcal{G}$ is termed homophilic if $\mathcal{H}_{\mbox{node}}(\mathcal{G})\geq 0.5$, and heterophilic otherwise.
Although homophily measures similarity across an entire graph, individual groups within a graph may display different homophily behaviors. Our ClassContrast approach leverages class interactions and introduces a class-aware homophily score through non-symmetric measures:
Definition 4.5 (ClassContrast (CC) Homophily Matrices).
@@ 4.4 Homophily and ClassContrast
In Appendix Table 9, we provide examples of CC homophily matrices. This matrix $\mathbf{M}$ provides detailed insights into both intra-class (homophily) and inter-class (heterophily) interactions, with diagonal elements indicating the propensity for nodes to connect within their own class.
@@
For a given $\alpha$ Homophily Matrix $\mathcal{M}^{\alpha}$, $\alpha$ Homophily $\mathcal{H}^{\alpha}_{cc}$ is defined as the average of the diagonal elements. i.e, $\mathcal{H}^{\alpha}_{cc}=\frac{1}{N}\sum_{i=1}^{N}h_{ii}$.
For example, the Spatial-1 Homophily ratio reveals a node’s likelihood to connect with nodes of the same class within its immediate neighborhood. Homophily matrices not only introduce new ways to measure homophily but also relates to various existing homophily metrics:
@@
For a given $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $\vec{\alpha}_{1}(v)$ be CC spatial-1 vector. Let $\widehat{\alpha}_{1}(v)$ be the vector where the entry corresponding to the class of $v$ is set to 00. Then, $1-\mathcal{H}_{\mbox{node}}(\mathcal{G})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{\|\widehat{\alpha}_{1}(v)\|_{1}}{\|\vec{\alpha}_{1}(v)\|_{1}}$.
The proof of this theorem and further discussions on how our ClassContrast homophily relates to other forms of homophily are detailed in the Appendix B.
@@ 5 Experiments
We evaluate ClassContrast in two tasks: node classification and link prediction. We share our Python implementation at https://anonymous.4open.science/r/CC-77A6/
@@
We use three widely-used homophilic datasets which are citation networks, CORA, CITESEER, and PUBMED (Sen et al. 2008), two Open Graph Benchmark (OGB) datasets: OGBN-ARXIV and OGBN-MAG (Hu et al. 2020), and five heterophilic datasets, TEXAS, CORNELL, WISCONSIN, CHAMELEON and SQUIRREL (Pei et al. 2019). The datasets are described in Appendix D and their statistics are given in Table 1.
Model Setup and Metrics.
For the three citation networks and five heterophilic datasets, the same setup is used as outlined by Bodnar et al. (Bodnar et al. 2022): we split nodes of each class into 48%, 32%, and 20% for training, validation and testing and report averages of 10 runs. We follow the train/validation/test split set by the OGB to report the accuracy score for OGBN-ARXIV and OGBN-MAG (Hu et al. 2020).
Embeddings.
We give the details of ClassContrast embeddings for each dataset in appendix D. The dimensions of the embeddings used for each dataset are given in Table 11.
Model Hyperparameters.
ClassContrast uses an MLP classifier (Haykin 1998). We configure the architecture with $700$ neurons for the hidden layers, a learning rate of $0.001$, and employed the Adam optimization method for training for 500 epochs. We utilized the ReLU activation function for the hidden layers and applied a L2 regularization value of $0.00001$ to mitigate overfitting.
Hardware.
We ran non-OGB experiments on a machine with Apple m2 chip Processor (8-core CPU, 10-core GPU, and, 6-core Neural Engine), and 16Gb of RAM. For the OGB data sets, we ran experiments on Google Colab with Intel(R) processor, 2.20GHz CPU, V100 GPUs, and 25.5Gb of RAM.
Runtime.
ClassContrast is computationally efficient. ClassContrast requires approximately $4$ hours for OGBN-ARXIV and $10$ hours for OGB-MAG to create all embeddings, and about $10$ minutes to train for all OGB datasets. End-to-end, our model processes three citation network datasets and three WebKb datasets, including the CHAMELEON dataset, in under a minute. Moreover, PUBMED takes about $20$ minutes and SQUIRREL approximately $3$ minutes. For comparison, RevGNN-Deep requires $13.5$ days and RevGNN-Wide takes $17.1$ days to train for 2000 epochs on a single NVIDIA V100 for the OGB datasets.
5.1 Node Classification Results
Baselines.
We compare ClassContrast with 12 state-of-the-art models (Table 2) and three classical models: GCN (Kipf and Welling 2017), GraphSage (Hamilton et al. 2017) and GAT (Veličković et al. 2018a). We choose ALT-APPNP (Liang et al. 2023) and LRGNN (Liang et al. 2023) SOTA models as they are recent and outperform the competitors in their respective papers. All baselines in Tables 2 and 3 use the transductive setting except for GraphSAGE, which uses the inductive setting.
@@ 5.1 Node Classification Results
[…]
@@
SQUIRREL 5,201 2,17,073 5 2,089 48/32/20 0.22
@@
Table 2: Node Classification Performance. Accuracy results for node classification. Baseline models from Geom-GCN to Gen-NSD were reported in (Bodnar et al. 2022; Li et al. 2022). Best results are given in bold, and the second best are underlined. The last column shows the average deviation of each model’s performance from the best performance across all datasets.
@@
Dataset CORA CITESEER PUBMED TEXAS CORNELL WISC. CHAM. SQUIR.
Node Homophily 0.83 0.72 0.79 0.10 0.39 0.15 0.25 0.22
GCN (Kipf and Welling 2017) 86.14$\pm$1.10 75.51$\pm$1.28 87.22$\pm$0.37 56.22$\pm$5.81 60.54$\pm$5.30 51.96 $\pm$5.17 65.94$\pm$3.23 49.63$\pm$2.43
GraphSAGE (Hamilton et al. 2017) 86.26$\pm$1.54 76.04$\pm$1.30 88.45$\pm$0.50 75.95$\pm$5.01 75.95$\pm$5.01 81.18$\pm$5.56 58.73$\pm$1.68 41.61$\pm$0.74
GAT (Veličković et al. 2018a) 85.03$\pm$1.61 76.55$\pm$1.23 87.30$\pm$1.10 54.32$\pm$6.30 61.89$\pm$5.05 49.41$\pm$4.09 60.26$\pm$2.50 40.72$\pm$1.55
Geom-GCN (Pei et al. 2019) 85.35$\pm$1.57 78.02$\pm$1.15 89.95$\pm$0.47 66.76$\pm$2.72 60.54$\pm$3.67 64.51$\pm$3.66 60.00$\pm$2.81 38.15$\pm$0.92
H2GCN (Zhu et al. 2020) 87.87$\pm$1.20 77.11$\pm$1.57 89.49$\pm$0.38 84.86$\pm$7.23 82.70$\pm$5.28 87.65$\pm$4.98 60.11$\pm$2.15 36.48$\pm$1.86
GPRGCN (Chien et al. 2020) 87.95$\pm$1.18 77.13$\pm$1.67 87.54$\pm$0.38 81.35$\pm$5.32 78.11$\pm$6.55 82.55$\pm$6.23 46.58$\pm$1.71 31.61$\pm$1.24
GCNII (Chen et al. 2020) 88.37$\pm$1.25 77.33$\pm$1.48 90.15$\pm$0.43 77.57$\pm$3.83 77.86$\pm$3.79 80.39$\pm$3.40 63.86$\pm$3.04 38.47$\pm$1.58
WRGAT (Suresh et al. 2021) 88.20$\pm$2.26 76.81$\pm$1.89 88.52$\pm$0.92 83.62$\pm$5.50 81.62$\pm$3.90 86.98$\pm$3.78 65.24$\pm$0.87 48.85$\pm$0.78
LINKX (Lim et al. 2021a) 84.64$\pm$1.13 73.19$\pm$0.99 87.86$\pm$0.77 74.60$\pm$8.37 77.84$\pm$5.81 75.49$\pm$5.72 68.42$\pm$1.38 61.81$\pm$1.80
NLGAT (Liu, Wang, and Ji 2021) 88.50$\pm$1.80 76.20$\pm$1.60 88.20$\pm$0.30 62.60$\pm$7.10 54.70$\pm$7.60 56.90$\pm$7.30 65.70$\pm$1.40 56.80$\pm$2.50
GloGNN++(Li et al. 2022) 88.33$\pm$1.09 77.22$\pm$1.78 89.24$\pm$0.39 84.05$\pm$4.90 85.95$\pm$5.10 88.04$\pm$3.22 71.21$\pm$1.84 57.88$\pm$1.76
GGCN (Yan et al. 2022) 87.95$\pm$1.05 77.14$\pm$1.45 89.15$\pm$0.37 84.86$\pm$4.55 85.68$\pm$6.63 86.86$\pm$3.29 71.14$\pm$1.84 55.17$\pm$1.58
Gen-NSD (Bodnar et al. 2022) 87.30$\pm$1.15 76.32$\pm$1.65 89.33$\pm$0.35 82.97$\pm$5.13 85.68$\pm$6.51 89.21$\pm$3.84 67.93$\pm$1.58 53.17$\pm$1.31
LRGNN (Liang et al. 2023) 88.30$\pm$0.90 77.50$\pm$1.30 90.20$\pm$0.60 90.30$\pm$4.50 86.50$\pm$5.70 88.20$\pm$3.50 79.16$\pm$2.05 74.38$\pm$1.96
ALT-APPNP (Xu et al. 2023) 89.60$\pm$1.30 79.90$\pm$1.20 90.30$\pm$0.50 89.50$\pm$2.20 90.40$\pm$4.50 88.60$\pm$3.30 77.00$\pm$1.90 69.40$\pm$1.50
ClassContrast 88.65$\pm$1.25 86.86$\pm$1.91 91.00$\pm$0.45 90.54$\pm$4.80 90.81$\pm$2.90 93.33$\pm$2.95 84.08$\pm$1.55 54.65$\pm$1.27
@@
Table 3: Classification accuracy of baselines and our ClassContrast model on OGBN datasets.
@@
GCN (Kipf and Welling 2017) 71.74 34.87
GraphSAGE (Hamilton et al. 2017) 71.49 37.04
GAT (Veličković et al. 2018a) 73.91 37.67
DeepGCN (Li et al. 2019) 72.32
DAGNN (Liu, Gao, and Ji 2020) 72.09
UniMP-v2 (Shi et al. 2020) 73.92
RevGAT (Li et al. 2021) 74.26
RGCN (Yu et al. 2022a) 47.96
HGT (Yu et al. 2022a) 49.21
R-HGNN (Yu et al. 2022a) 52.04
LEGNN (Yu et al. 2022b) 73.71 53.78
ClassContrast 71.56 50.03
Performance.
We present the classification results in Table 2 and Table 3. OGB results are shown in Table 3. ClassContrast performance is within 3% of the SOTA accuracy for the OGB datasets. Table 2 shows that ClassContrast outperforms SOTA models in 6 out of 8 datasets. In particular, ClassContrast achieves a median accuracy of $89.60$ compared to $89.05$ of the best baseline, ALT-APPNP (Xu et al. 2023). The improvements are most notable in the CITESEER ($+6.96$) and Chameleon ($+7.08$) datasets, underlining the contrasive power of our approach in both homophilic and heterophilic datasets.
SQUIRREL performance.
An exception to the SOTA values in Table 2 is the SQUIRREL dataset, which allows us to showcase the use of energy landscape theory through homophily, as follows.
SQUIRREL homophily matrices in  Table 6 reveal significant diversity in class interactions, particularly evident in the "Spatial-1" and "Spatial-2" matrices. These matrices show substantial inter-class interactions, sometimes exceeding intra-class interactions. Such a pattern can be interpreted using energy landscape theory, where the energy barriers between different classes are low, suggesting a disordered energy landscape. This scenario indicates that the energy required for a node to transition from one class to another, in terms of classification, is reduced. Consequently, there isn’t a well-defined ’valley’ or ’well’ that confines nodes strictly within one class, complicating the classification model’s ability to accurately define class boundaries and leading to lower accuracy. Please see  Table 10 where the inter-calls interactions are more moderate for the Wisconsin dataset. Furthermore, we can contrast the SQUIRREL and WISCONSIN matrices with those of the highly homophilous CORA dataset in Table 9, where diagonal values are more pronounced.
Domain homophily matrix in Table 6 shows some improvement in intra-class interactions, yet the highest homophily ratios remain relatively modest (around $0.22-0.23$). This implies that while domain features do provide some level of class-specific clustering, the low energy states are shallow, allowing nodes to exhibit characteristics similar to multiple classes. Such shallow states suggest that during the learning process, nodes are not distinctly categorized into deep, well-separated classes, which can influence the training dynamics and lead to blurred class separations.
GNNs with ClassContrast Embeddings.
To evaluate the integration of ClassContrast (CC) with GNNs, we replace the initial node embeddings in GNNs with CC embeddings and assess their effectiveness. Accuracy results are presented in Table 4 and Figures 3 and 4 with additional details given in Appendix A. The results indicate that CC embeddings accelerate the convergence of GNN performance while maintaining a consistent accuracy gap over the vanilla model. Furthermore, combining CC with the recent LINKX model yields the best results across all baselines for both datasets, highlighting a promising avenue for enhancing GNN performance.
Table 4: Vanilla-GNN vs. GNN+CC accuracy results for node classification. See Appendix A for details.
Dataset Model GNN CC-GNN Imp.($\uparrow$)
CORA GCN 86.14$\pm$1.10 88.43$\pm$0.92 2.29
SAGE 86.26$\pm$1.54 90.08$\pm$1.37 3.82
GAT 85.03$\pm$1.61 87.18$\pm$2.12 2.15
LINKX 84.64$\pm$1.13 90.06$\pm$1.28 5.42
TEXAS GCN 56.22$\pm$5.81 70.81$\pm$6.43 14.59
SAGE 75.95$\pm$5.01 87.84$\pm$6.65 11.89
GAT 54.32$\pm$6.30 62.16$\pm$5.70 7.84
LINKX 74.60$\pm$8.37 93.78$\pm$4.04 19.18
Figure 3: Performance comparison of three GNN models (GCN, GSAGE, GAT) starting with the original node embeddings (blue) and ClassContrast node embeddings (orange) on CORA dataset.
@@
Table 5: Link Prediction Performances. AUC results for Link Prediction. Baselines are reported from (Fu et al. 2023; Zhou et al. 2022). In the Overall column, we report mean AUC results across all datasets.

### v2

@@ 3.3 SCNode Embedding
Definition 3.4 (SCNode Embedding).
The final embedding $\vec{\gamma}(u)$ for node $u$ is a concatenated vector of these embeddings as shown in the right. For example, if the graph is directed, and we want to utilize two landmark types, $\{\xi^{1}_{j}\}$ and $\{\xi^{2}_{j}\}$, the resulting SCNode vector will be in a $5\cdot N$-dimensional vector where $N$ is the number of classes. For clarity, we represent $\vec{\gamma}(u)$ in a 2D format ($m\times N$), where each column corresponds to one class and each row represents one type of SCNode vector.
$\vec{\gamma}(u)=\left[\begin{array}[]{lcr}\longleftarrow&\vec{\alpha}_{1i}(u)&\longrightarrow\\ \longleftarrow&\vec{\alpha}_{1o}(u)&\longrightarrow\\ \longleftarrow&\vec{\alpha}_{2}(u)&\longrightarrow\\ \longleftarrow&\vec{\beta}_{1}(u)&\longrightarrow\\ \longleftarrow&\vec{\beta}_{2}(u)&\longrightarrow\end{array}\right]\begin{array}[]{ll}\text{Spatial}&\text{Incoming 1-ngbd}\\ \text{Spatial}&\text{Outgoing 1-ngbd}\\ \text{Spatial}&\text{2-ngbd}\\ \text{Contextual}&\{d(\mathbf{X}_{u},\xi^{1}_{j})\}\\ \text{Contextual}&\{d(\mathbf{X}_{u},\xi^{2}_{j})\}\end{array}$
Relation to label-structure methods.
SCNode augments each node with two complementary signals: contextual coordinates given by distances to class landmarks in the embedding space, and spatial class-distribution histograms over k-hop neighborhoods. This encodes label structure over nodes without modifying the graph or introducing virtual nodes. In contrast, techniques such as LEGNN (Yu et al., 2022b) and GraphGPS (Rampášek et al., 2022) inject graph-level signals or virtual nodes derived from class labels. Conceptually, the class landmarks play the role of prototypes: distances to class-conditioned centroids provide barycentric coordinates in label space (Snell et al., 2017), but unlike prototypical networks (Hou & Sato, 2022), our prototypes are computed from the training graph and kept fixed during message passing, which decouples metric learning from the backbone and enables transductive inference. Our approach is backbone agnostic and operates at the feature level, so it pairs naturally with standard GNNs and graph transformers.
3.4 Homophily and SCNode
In this section, we demonstrate the effectiveness of the SCNode framework in analyzing homophily behavior within the graph from multiple perspectives and gaining deeper insights into class interactions. In recent years, several metrics have been introduced to study the effect of homophily on graph representation learning (Lim et al., 2021b; Jin et al., 2022; Luan et al., 2022; Platonov et al., 2023a) (see overview in Section˜A.1). A widely used metric, the node homophily ratio, is defined as $\mathcal{H}_{\mbox{node}}(\mathcal{G})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{\eta(v)}{deg(v)}$, where $\eta(v)$ represents the number of adjacent nodes to $v$ sharing the same class. A graph $\mathcal{G}$ is termed homophilic if $\mathcal{H}_{\mbox{node}}(\mathcal{G})\geq 0.5$, and heterophilic otherwise.
Although homophily measures similarity across an entire graph, individual groups within a graph may display different homophily behaviors. Our SCNode approach leverages class interactions and introduces a class-aware homophily score through non-symmetric measures:
Definition 3.5 (SCNode Homophily Matrices).
@@ 3.4 Homophily and SCNode
Table˜9 in the Appendix presents examples of SCN homophily matrices. These matrices offer detailed insights into intra-class (homophily) and inter-class (heterophily) interactions across spatial and contextual contexts, with the diagonal elements representing the likelihood of nodes connecting within their own class.
@@
For a given $\alpha$ Homophily Matrix $\mathcal{M}^{\alpha}$, $\alpha$ Homophily $\mathcal{H}^{\alpha}_{scn}$ is defined as the average of the diagonal elements. i.e, $\mathcal{H}^{\alpha}_{scn}=\frac{1}{N}\sum_{i=1}^{N}h_{ii}$.
For example, the Spatial-1 Homophily ratio reveals a node’s likelihood to connect with nodes of the same class within its immediate neighborhood. Homophily matrices not only introduce new ways to measure homophily but also relate to various existing homophily metrics:
@@
For a given $\mathcal{G}=(\mathcal{V},\mathbb{E})$, let $\vec{\alpha}_{1}(v)$ be SCN spatial-1 vector. Let $\widehat{\alpha}_{1}(v)$ be the vector where the entry corresponding to the class of $v$ is set to 0. Then, $1-\mathcal{H}_{\mbox{node}}(\mathcal{G})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{\|\widehat{\alpha}_{1}(v)\|_{1}}{\|\vec{\alpha}_{1}(v)\|_{1}}$.
The proof of the theorem and further discussions on how our SCNode homophily matrices relate to other forms of homophily metrics are detailed in the Appendix˜A.
@@ 4 Experiments
We evaluate SCNode in two tasks: node classification and link prediction. We share our Python implementation at https://github.com/joshem163/SCNode
@@
We use three widely-used homophilic datasets, which are citation networks, CORA, CITESEER, and PUBMED (Sen et al., 2008), two Open Graph Benchmark (OGB) datasets: OGBN-ARXIV and OGBN-MAG (Hu et al., 2020), and ten heterophilic datasets, including TEXAS, CORNELL, WISCONSIN, and CHAMELEON (Pei et al., 2019), as well as filtered versions of SQUIRREL and CHAMELEON, AMAZON-RATINGS, TOLOKER, and QUESTIONS (Platonov et al., 2023b). The datasets are described in Appendix˜C and their statistics are given in Table˜1.
@@
Datasets Nodes Edges Classes Features Tr/Val/Test ($\%$) Hom.
@@
SQUIRREL 2,223 46,998 5 2,089 48/32/20 0.19
CHAMELEON 890 8,854 5 2,325 48/32/20 0.24
ROMAN-EMPIRE 22,662 32,927 18 300 50/25/25 0.05
AMAZON-RATINGS 24,492 93,050 5 300 50/25/25 0.37
TOLOKERS 11,758 5,19,000 2 10 50/25/25 0.63
QUESTIONS 48,921 153,540 2 301 50/25/25 0.89
denotes the filtered version of the dataset.
Experimental Setup.
We evaluate SCNode on three citation networks and four heterophilic graph datasets, following the experimental protocol established by Bodnar et al. (Bodnar et al., 2022). For these datasets, we randomly split the nodes of each class into $48\%$ for training, $32\%$ for validation, and $20\%$ for testing, and report the average accuracy over 10 independent runs. For the OGB datasets, OGBN-ARXIV and OGBN-MAG, we use the official train/validation/test splits provided by the Open Graph Benchmark (OGB) (Hu et al., 2020), and report the classification accuracy accordingly. For the remaining datasets, we adopt the same splitting strategy as outlined in (Platonov et al., 2023b). We give the details of SCNode embeddings for each dataset in  Appendix˜C. The dimensions of the embeddings used for each dataset are given in Table˜11.
For classification, we employ a Multi-Layer Perceptron (MLP) with a single hidden layer consisting of 100 neurons. We use the ReLU activation function, a learning rate of 0.001, and the Adam optimizer for training up to 500 epochs. To mitigate overfitting, we apply L2 regularization with a weight decay of 0.0001.
All non-OGB experiments were conducted on a local machine equipped with an Apple M2 chip (8-core CPU, 10-core GPU, 6-core Neural Engine) and 16 GB of RAM. OGB experiments were run on Google Colab, using a system with an Intel(R) 2.20GHz CPU, NVIDIA V100 GPU, and 25.5 GB of RAM.
Runtime.
SCNode is computationally efficient. SCNode requires approximately $4$ hours for OGBN-ARXIV and $10$ hours for OGB-MAG to create all embeddings, and about $10$ minutes to train for all OGB datasets. End-to-end, our model processes three citation network datasets and three WebKb datasets, including the CHAMELEON dataset, in under a minute. PUBMED takes about $20$ minutes. For comparison, RevGNN-Deep requires $13.5$ days and RevGNN-Wide takes $17.1$ days to train for 2000 epochs on a single NVIDIA V100 for the OGB datasets (Li et al., 2021).
4.1 Node Classification Results
Baselines.
We compare SCNode against a range of state-of-the-art models( Tables˜2, 3 and 5), including three classical approaches: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018a), as well as a graph transformer model. All baselines in Tables˜2, 3 and 5 are evaluated in the transductive setting, except for GraphSAGE, which follows the inductive setting.
Performance.
We present the node classification results for benchmark homophilic and heterophilic datasets in Tables˜2 and 3 and for OGB datasets in Table˜5. On the two homophilic benchmarks, SCNODE achieves state-of-the-art performance, consistently outperforming prior methods. Even more strikingly, on the four heterophilic graphs, SCNODE delivers dramatic gains, surpassing the previous best by 2 – 4 points. For the remaining five heterophilic datasets, SCNODE achieves the best results on two and demonstrates competitive performance on the others, as shown in Table˜3. This uniform superiority across low- and high-homophily settings underscores SCNODE’s ability to adaptively integrate both local feature similarity and global topological cues, yielding a single architecture that excels regardless of underlying graph structure. Moreover, on the OGB benchmarks, SCNODE achieves results within 3-4 % of the state-of-the-art, despite relying on a far more compact architecture than competing deep-learning models.
@@ 4.1 Node Classification Results
Table 2: Node Classification Performance. Accuracy results for node classification tasks. Baseline results up to Gen-NSD are sourced from (Bodnar et al., 2022; Li et al., 2022), while the remaining results are taken from their respective original papers. The best results are shown in bold, while the second-best results are underlined. The last column reports the average performance gap (%) relative to the best result across all datasets.
@@
Dataset CORA CITESEER PUBMED TEXAS CORNELL WISC. CHAM. Avg. Gap
Node Homophily 0.83 0.72 0.79 0.10 0.39 0.15 0.25 %
[…]
Table 3: More Heterophilic Datasets. Node classification results of our models compared to GNN and graph transformer baselines on heterophilic benchmarks. The last column reports the average gap to the best-performing model in each dataset. Baseline models are sourced from (Platonov et al., 2023b; Yadati, 2025; Luo et al., 2024b)
Model Squirrel* Chameleon* Amazon Tolokers Questions Avg. Gap
[…]
@@
Table 4: Classification accuracy of baselines and our SCNode model on OGBN datasets.
@@
GCN (Kipf & Welling, 2017) 71.74 34.87
GSAGE (Hamilton et al., 2017) 71.49 37.04
GAT (Veličković et al., 2018a) 73.91 37.67
DeepGCN (Li et al., 2019) 72.32
DAGNN (Liu et al., 2020) 72.09
UniMP-v2 (Shi et al., 2020) 73.92
RevGAT (Li et al., 2021) 74.26
RGCN (Yu et al., 2022a) 47.96
HGT (Yu et al., 2022a) 49.21
R-HGNN (Yu et al., 2022a) 52.04
LEGNN (Yu et al., 2022b) 73.71 53.78
S3GCL (Wan et al., 2024) 71.36
LMSPS (Liu et al., 2024) 54.83
SCNode 71.56 50.03
Table 5: Vanilla-GNN vs. SCNode+GNN (SCN-GNN) accuracy results for node classification.
Dataset Model GNN SCN-GNN Imp.($\uparrow$)
[…]
SCN-GNNs: GNNs with SCNode Embeddings. To evaluate the integration of SCNode with GNNs as plug-and-play components, we replaced the initial node embeddings in GNNs with SCN embeddings and assessed their effectiveness. We tested three classical GNN models (GCN, GraphSAGE, and GAT), along with the recent LINKX model. A two-layer GNN framework was implemented using the Adam optimizer with an initial learning rate of 0.001, a dropout rate of 0.5, a weight decay of 5e-4, and 32 hidden channels. The results, shown in Table˜5 and Figures˜3 and 4, indicate significant performance improvements when using SCN embeddings. These embeddings accelerate convergence and maintain a consistent accuracy advantage over the vanilla models. Notably, combining SCNode with the LINKX model yields the best results across all baselines for both datasets, showcasing the potential of SCNode to enhance GNN performance. This improvement is attributed to the effective integration of spatial and contextual information provided by SCNode embeddings. The accuracy gap remains stable, further demonstrating their robustness. In future work, we aim to explore this integration more comprehensively to further enhance GNN performance.
@@
Table 6: Link Prediction Performances. AUC results for Link Prediction. Baselines are reported from (Fu et al., 2023; Zhou et al., 2022; Wu et al., 2024). In the Overall column, we report mean AUC results across all datasets.